OpenStudy (anonymous):

find x intercept of a parabola when vertex is (1,1) and y-intercept is (0,-3)

OpenStudy (anonymous):

ans: x -intercepts at 1/2 and 3/2

OpenStudy (anonymous):

Oh it says y oh ok

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first we can find the equation for the parabola then we can set it equal to zero and solve

OpenStudy (misty1212):

in vertex from is looks like \[y=a(x-1)^2+1\] and we need \(a\)

OpenStudy (misty1212):

if \(x=0\) you are told that \(y=-3\) then we get \[-3=a(0-1)^2+1\]or \[-3=a+1\] making \[a=-4\]

OpenStudy (misty1212):

set \[-4(x-1)^2+1=0\] and solve for \(x\)